OpenStudy (anonymous):

[9.06] Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 2, – 16), and x intercepts at x = -6 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 4x − ___

OpenStudy (anonymous):

one again im back msellier sister need help with that please? @mathslover @ivettef365 @shkrina @.Sam. @Jamierox4ev3r

terenzreignz (terenzreignz):

Well, if you have x-intercepts at points x = a and x = b the most natural quadratic equation to check first would be y = (x - a)(x - b)

terenzreignz (terenzreignz):

In this case, y = (x + 6)(x - 2) Solve that, and the missing term will be clear :)

OpenStudy (anonymous):

?

terenzreignz (terenzreignz):

Evaluate (x + 6)(x - 2) using the FOIL method, or any method you've learned on how to multiply binomials...

OpenStudy (anonymous):

well thats the whole of me asking i have trouble with that ?

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Multiplying binomials using the FOIL method?

terenzreignz (terenzreignz):

The FOIL method is one of the more basic methods of multiplying a pair of binomials. FOIL is actually an acronym. Let's demonstrate with a generic pair of binomials: \[\Large (a+b)(p+q)\] Now, let's break down the acronym FOIL into its individual letters: FIRST terms: \[\Large (\color{red}a+b)(\color{red}p+q)\]\[\Large (a+b)(p+q)=\color{red}{ap}...\](the ... indicates we're not done yet) Now: OUTER terms: \[\Large (\color{green}a+b)(p+\color{green}q)=ap + \color{green}{aq}...\] Then: INNER terms: \[\Large (a+\color{orange}b)(\color{orange}p+q)=ap+aq + \color{orange}{bp}...\] Finally: LAST terms: \[\Large (a+\color{blue}b)(p+\color{blue}q)=ap+aq+bp+\color{blue}{bq}\] So, as per the FOIL METHOD (first, outer, inner, last), the product of these two binomials is simply \[\Large (a+b)(p+q)=\color{red}{ap}+\color{green}{aq}+\color{orange}{bp}+\color{blue}{bq}\]

OpenStudy (anonymous):

the things is i need help of how to solve im bad at this

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

(x + 6)(x - 2) First terms, what are they?

OpenStudy (anonymous):

x=ay+by/d solve for y

OpenStudy (anonymous):

please

OpenStudy (anonymous):

the answer is